OpenStudy (anonymous):

solve for R ; rx+9/5 = h

pooja195 (pooja195):

Hi :-) Welcome to openstudy ^_^ Any ideas on where to start? Have you attempted the problem yet?

OpenStudy (anonymous):

I've attempted but not getting any where

pooja195 (pooja195):

Alright show me what you have so far :)

OpenStudy (anonymous):

rx+9/5=h -9 -9 ------------ rx/5=h-9 --- = h-9/x x r/5=h-9/x *5=h-9/x*5 r=h-9/x*5

pooja195 (pooja195):

Hmm :-/ @AlexandervonHumboldt2

OpenStudy (anonymous):

can you help @AlexandervonHumboldt2 and @pooja195

pooja195 (pooja195):

@♪Chibiterasu